Find the length of a rectangular lot with a perimeter of 92m if the length is 6m more than the width
Answer by lwsshak3(11628)

Find the length of a rectangular lot with a perimeter of 92m if the length is 6m more than the width
let x=width
x+6=length
2(width+length=perimeter
2(x+x+6)=92
4x+12=92
4x=80
x=20
x+6=26
width =20 m
length=26 m